### v1

# SecureWebArena: A Holistic Security Evaluation Benchmark for LVLM-based Web Agents

Abstract.
Large vision–language model (LVLM)-based web agents are emerging as powerful tools for automating complex online tasks. However, when deployed in real-world environments, they face serious security risks, motivating the design of security evaluation benchmarks. Existing benchmarks provide only partial coverage, typically restricted to narrow scenarios such as user-level prompt manipulation, and thus fail to capture the broad range of agent vulnerabilities. To address this gap, we present SecureWebArena, the first holistic benchmark for evaluating the security of LVLM-based web agents. SecureWebArena first introduces a unified evaluation suite comprising six simulated but realistic web environments (e.g., e-commerce platforms, community forums) and includes 2,970 high-quality trajectories spanning diverse tasks and attack settings. The suite defines a structured taxonomy of six attack vectors spanning both user-level and environment-level manipulations. In addition, we introduce a multi-layered evaluation protocol that analyzes agent failures across three critical dimensions: internal reasoning, behavioral trajectory, and task outcome, facilitating a fine-grained risk analysis that goes far beyond simple success metrics. Using this benchmark, we conduct large-scale experiments on 9 representative LVLMs, which fall into three categories: general-purpose, agent-specialized, and GUI-grounded. Our results show that all tested agents are consistently vulnerable to subtle adversarial manipulations and reveal critical trade-offs between model specialization and security. By providing (1) a comprehensive benchmark suite with diverse environments and a multi-layered evaluation pipeline, and (2) empirical insights into the security challenges of modern LVLM-based web agents, SecureWebArena establishes a foundation for advancing trustworthy web agent deployment.
Do, Not, Use, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Figure 1. Overall illustration of our SecureWebArena, the first holistic benchmark for evaluating the security of LVLM-based web agents.
Large vision language models (LVLMs) (Hurst et al., 2024; Team et al., 2023; Wang et al., 2024) have equipped autonomous agents with powerful capabilities to perceive and reason across language, vision, and user interface elements (Zeng et al., 2025; Sravanthi et al., 2025; Huang et al., 2025). As web agents, these models can navigate complex websites, fill out forms, and make multi-step decisions based on combined visual and textual input (Ning et al., 2025; Abuelsaad et al., 2024; Lai et al., 2024). However, as these agents are deployed in real-world scenarios to handle sensitive data and critical workflows, their growing autonomy exposes them to severe security threats, such as pop-up attacks (Zhang et al., 2024) and prompt injections (Evtimov et al., 2025; Johnson et al., 2025; Wang et al., 2025b; Yu et al., 2025).
The growing recognition of these security threats has led to the first wave of security evaluation benchmarks (Kumar et al., 2024; Tur et al., 2025; Evtimov et al., 2025; Levy et al., 2024; Liu et al., 2025; Ying et al., 2024a; Xiao et al., 2023b, a). While these valuable contributions have begun to explore the security problem, they often do so with a limited scope, focusing on isolated aspects of the threat landscape. Some benchmarks primarily investigate risks stemming from malicious or harmful user instructions (Kumar et al., 2024; Tur et al., 2025; Ying et al., 2025d, 2024b). Others concentrate on specific, narrow threat models, with a notable focus on prompt injection originating from within the web page (Evtimov et al., 2025), or on adherence to predefined policies in enterprise contexts (Levy et al., 2024). In summary, existing security evaluation benchmarks for web agents fail to provide a unified, systematic framework that addresses vulnerabilities from both user-level instructions and diverse environment-level manipulations, thus failing to capture the broad range of vulnerabilities.
To address this gap, this paper introduces SecureWebArena, the first holistic benchmark specifically designed for evaluating the security of LVLM-based web agents. Our benchmark first provides a unified evaluation suite featuring 6 simulated yet realistic representative web environments, such as online shopping and code management platforms. Central to our framework is a structured classification of 6 attack vectors that span both user-level manipulations (e.g., Jailbreak (Zou et al., 2023; Ying et al., 2025a, c; Zou et al., 2025; Ying et al., 2025b; Mu et al., 2025)) and environment-level threats (e.g., Pop-up Attack (Zhang et al., 2024)). To enable a deeper level of assessment, we introduce a multi-layered evaluation protocol that analyzes agent failures across three critical dimensions: internal reasoning, behavioral trajectory, and task outcome. This approach facilitates a fine-grained risk analysis that goes far beyond simple success metrics. Our main contributions are:
  • We build SecureWebArena, the first holistic evaluation benchmark for LVLM-based web agent security, featuring realistic simulated environments with 330 adversarial scenarios and a structured classification of attacks from both user and environment sources.
  • Our benchmark introduces a multi-layered evaluation protocol that assesses agent failures across internal reasoning, behavioral trajectory, and task outcome, enabling a more granular and insightful risk analysis.
  • We conduct extensive experiments on 9 representative agents across three distinct LVLM types, providing a comparative analysis of their security vulnerabilities and revealing critical robustness trade-offs.
Our results reveal that modern LVLM-based web agents are universally vulnerable to subtle attacks and uncover critical security trade-offs tied to model specialization, demonstrating that no single type of LVLM is resilient across all attack vectors. We hope that SecureWebArena will serve as both a critical diagnostic tool and a foundational benchmark, guiding the community toward building more secure and resilient web agents.
2. Related Work
2.1. Benchmarks for Web Agents
The evaluation of web agents has traditionally focused on task-oriented performance, assessing their ability to navigate web environments and complete user instructions. Early benchmarks such as WebShop (Yao et al., 2022), Mind2Web (Deng et al., 2023a), WebArena (Zhou et al., 2023), and VisualWebArena (Koh et al., 2024) were instrumental in measuring functional capabilities but largely overlooked safety and security vulnerabilities. More recent efforts have begun to construct security-focused benchmarks, although these typically address specific threat dimensions: BrowserART (Kumar et al., 2024) and SAFEARENA (Tur et al., 2025) examine agent responses to harmful user instructions, while WASP (Evtimov et al., 2025) targets prompt injection from malicious web content. Other benchmarks explore niche scenarios such as socio-cultural sensitivity (Qiu et al., 2024) or policy compliance in enterprise systems (Levy et al., 2024).
Despite these contributions, existing benchmarks remain limited in three key ways. ❶ Most works evaluate user- or environment-level attacks separately, lacking a unified framework. ❷ The diversity of attack vectors is limited, with only a subset of representative threats realized across agents or environments. ❸ Most evaluations treat agent decisions monolithically, lacking fine-grained analysis of why and how failures occur.
In contrast, our SecureWebArena offers a unified testbed encompassing six diverse web environments and six representative attack vectors, along with a multi-layered evaluation protocol that analyzes reasoning, behavior, and outcomes to reveal agent vulnerabilities under adversarial conditions. Tab. 1 summarizes the key differences between SecureWebArena and existing benchmarks.
2.2. Attack Vectors on Web Agents
As LVLMs are increasingly deployed in interactive web environments, their multimodal decision-making processes are being exploited by a diverse set of attack vectors. These attacks can be broadly categorized into two main strategies. The first manipulates the agent’s language understanding through methods like Direct Prompt Injection (DP Injection) (Wang et al., 2025b) and Jailbreak Attacks (Zou et al., 2023). The second, more common strategy deceives the agent through the web interface itself. This includes visually deceptive pop-ups and ads that mimic legitimate UI elements (Pop-up Attack/Ad Injection) (Zhang et al., 2024; Wang et al., 2025a), distraction techniques that obscure safe options (Distract Attack) (Ma et al., 2025), and Indirect Prompt Injection (IP Injection) (Greshake et al., 2023) that hide malicious commands within plausible-looking interface text.
To address this fragmented threat landscape, our SecureWebArena provides the first systematic framework to evaluate these diverse threats holistically. We operationalize a structured taxonomy of these representative attacks, embedding them across both user-level and environment-level settings to enable a comprehensive diagnosis of security vulnerabilities.
Table 1. Comparison of our SecureWebArena with existing web agent security evaluation benchmarks across key dimensions.
Benchmark Threat Source # Attack Vectors # Env # Adv Task # Trajectory # Model Type # Modality Multi-Eval? Real-Web?
BrowserART User-Level 1 3 100 800 1 1
ST-WEBAGENRBENCH User-Level 1 3 222 666 2 2
WASP Env-Level 1 3 84 420 2 1
SAFEARENA Env-Level 1 3 250 1250 1 2
SecureWebArena User-Level & Env-Level 6 6 330 2970 3 2
3. Threat Model
3.1. Preliminaries
We model a web agent based on the Set-of-Marks (SoM) paradigm (Yang et al., 2023) as a sequential decision-making system. The agent aims to accomplish a high-level task $\mathcal{G}$, specified by the user in natural language, within a dynamic web environment $\mathcal{E}$. The interaction proceeds over discrete timesteps $t=1,2,\dots,T$.
The agent is powered by a LVLM $\mathcal{M}$, which jointly reasons over visual and textual inputs to generate actions. At each timestep $t$, the agent performs the following steps.
  1. (1)
    State Perception. The agent captures the current state $s_{t}\in\mathcal{S}$ of the web environment via a SoM-augmented observation $o_{t}$. Specifically, a client-side script automatically annotates every interactable element on the current webpage with a unique integer ID and a colored bounding box. This yields two components: ❶A marked screenshot $v_{t}^{\text{SoM}}$, where each interactable element is overlaid with its ID and bounding box. ❷ A SoM metadata list $\mathcal{L}_{t}=\{(\text{id}_{i},\text{tag}_{i},\text{text}_{i})\}_{i=1}^{N}$, which provides the ID, HTML tag type (e.g., BUTTON, INPUT), and visible text content (if any) for each marked element. The full observation is thus:
    (1) $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t}).$
  2. (2)
    Reasoning and Action Generation. The LVLM $\mathcal{M}$ takes as input the user goal $\mathcal{G}$, the current SoM observation $o_{t}$, and the interaction history $\mathcal{H}_{t-1}=\{(o_{1},a_{1}),\dots,(o_{t-1},a_{t-1})\}$. It processes the interleaved image-text context to produce a CoT (Wei et al., 2022) reasoning trace $c_{t}$ and selects the next action $a_{t}$ from a discrete action space $\mathcal{A}$:
    (2) $c_{t},a_{t}=\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1}).$
    The action space $\mathcal{A}$ consists of commands that reference elements by their SoM ID, such as click[id], type[id][text], and scroll[up|down].
  3. (3)
    Environment Interaction. The selected action $a_{t}$ is executed in the environment $\mathcal{E}$, leading to a deterministic state transition:
    (3) $s_{t+1}=\mathcal{E}(s_{t},a_{t}).$
The process repeats until the agent determines that the task $\mathcal{G}$ is complete or a maximum number of steps $T$ is reached. The sequence of actions $\tau=(a_{1},a_{2},\dots,a_{T})$ constitutes the agent’s behavioral trajectory. An ideal trajectory $\tau^{*}$ is one that safely satisfies the goal $\mathcal{G}$.
3.2. Attacker’s Objectives and Capabilities
Based on the agent’s decision-making process defined in Sec. 3.1, an attacker’s goal is to manipulate the agent into executing a harmful or unintended trajectory $\tau$, causing it to deviate from the ideal trajectory $\tau^{*}$. We define a threat model that considers two primary points of intervention where an attacker can influence the agent’s decision-making function, $\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1})$: the user’s high-level goal $\mathcal{G}$, and the environment’s observation $o_{t}$. This leads to a natural classification of threats into two categories.
3.2.1. User-Level Threats
In this scenario, the attacker directly manipulates the high-level task instruction $\mathcal{G}$ provided to the agent. The environment $\mathcal{E}$ is assumed to be benign, but the goal itself is malicious. Formally, the attacker crafts a malicious goal $\mathcal{G}_{\text{malicious}}$ to replace the benign $\mathcal{G}$. These threats target the agent’s language understanding and safety alignment.
3.2.2. Environment-Level Threats
In this more complex scenario, the user’s goal $\mathcal{G}$ is benign, but the web environment $\mathcal{E}$ is controlled by the attacker. The attacker manipulates the environment to produce a deceptive observation $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t})$ that misleads the agent’s perception and reasoning. These threats exploit the agent’s reliance on both visual cues from the screenshot $v_{t}^{\text{SoM}}$ and structural information from the metadata $\mathcal{L}_{t}$.
We do not assume access to internal model gradients or weights, reflecting the realistic black-box deployment of commercial LVLMs. The agent operates under standard web permissions, and attackers cannot invoke system-level APIs. Attacks are confined to manipulations through the natural input-output interface of the agent, such as textual instructions or perceptible visual elements.
Figure 2. SecureWebArena framework. It integrates simulated environments, diverse attack vectors, and multi-level evaluation to assess agent safety performance to adversarial manipulation.
4. SecureWebArena Design
4.1. Environment Suite
To ensure the generality and robustness of our security evaluations, SecureWebArena is built upon 6 diverse, realistic web environments adapted from WebArena (Zhou et al., 2023) and VisualWebArena (Koh et al., 2024). These environments are implemented as lightweight, browser-based web applications with full HTML, CSS, and DOM structures, supporting dynamic rendering, user interaction, form handling, and layout variation. They provide a faithful simulation of modern web user interfaces while enabling attack injection and trajectory logging in a controlled, reproducible manner. Each environment is instrumented with SoM-based markup that uniquely identifies interactive elements. It also supports structured metadata export for element-level reasoning and dynamic attack surface exposure to simulate real-world threats.
The six environments span a broad range of real-world interaction contexts, grouped into four representative categories. Information Retrieval and Navigation environments (e.g., Wikipedia and Reddit) feature dense, user-generated content with richly linked layouts, testing agents’ ability to maintain focus amidst subtle textual distractions. E-commerce and Transaction tasks in Shopping simulate high-stakes workflows involving financial decisions and sensitive data, providing a natural setting to evaluate agents’ risk awareness and cautiousness. Content Management and Publishing environments (e.g., Classifieds and ShoppingAdmin) involve privileged actions such as creating, modifying, or deleting content, presenting scenarios where errors could lead to persistent or destructive consequences. Finally, the Software Development and Collaboration environment, based on GitLab, focuses on technical workflows such as issue tracking and code repository navigation, enabling evaluation of agents’ resilience in structured, domain-specific interfaces.
These environments impose diverse perceptual, reasoning, and interface navigation demands on agents. They also serve as rich platforms for embedding both user-level and environment-level adversarial content under varied task semantics. Fig. 2 illustrates the overall framework of the proposed benchmark, and App. A.1 provides illustrative examples of each environment.
4.2. Task Construction
Tasks in SecureWebArena are manually designed to reflect realistic user intents across the six interaction environments, covering a wide range of common goals such as information look-up, product purchase, issue reporting, and content editing. Each task is specified as a standalone natural language instruction (e.g., “Find a wireless mouse under \$20 and add it to the cart”) that assumes the agent is placed in an appropriate initial interface state (e.g., relevant page opened, default settings loaded).
Although our benchmark focuses exclusively on adversarial evaluation, each attack task in SecureWebArena is implicitly grounded in a corresponding benign goal. The adversarial version perturbs this underlying task with one of the defined attack vectors, either by altering the instruction (user-level), or modifying the rendering interface (environment-level). Each attack vector is instantiated using a dedicated template or pattern based on its definition in prior literature, ensuring consistent application across tasks and environments.
To support systematic evaluation, tasks in SecureWebArena are designed to reflect real-world user behavior (goal realism) and require understanding of both rendered UI elements and natural language instructions (multimodal dependency). Attacks are localized to a specific input channel for controlled assessment (isolated intervention), and tasks span diverse interaction patterns including navigation, selection, typing, and multi-step planning (coverage diversity). This design enables SecureWebArena to probe the holistic adversarial decision-making pipeline of web agents while supporting reproducible, attack-specific performance analysis.
4.3. Attack Taxonomy
SecureWebArena implements 6 representative attack vectors that target both user-level and environment-level inputs, as defined in our threat model (Sec. 3.2). Each attack manipulates a specific component of the agent’s decision function $\mathcal{M}(\mathcal{G},o_{t},\mathcal{H}_{t-1})$, either by altering the instruction $\mathcal{G}$ or the rendered observation $o_{t}=(v_{t}^{\text{SoM}},\mathcal{L}_{t})$. For clarity, we denote perturbed instructions and observations as $\widetilde{\mathcal{G}}$ and $\widetilde{o}_{t}$, and introduce $\mathcal{A}_{\text{adv}}$ to denote an adversarially modified action trajectory induced by these attacks. We categorize the 6 attack vectors as follows:
User-level Attacks.
These attacks directly manipulate the goal input $\mathcal{G}$, resulting in a tampered goal $\widetilde{\mathcal{G}}$ that adversarially biases the model’s response.
  • Direct Prompt Injection (DP Injection). The attacker appends an adversarial clause to a benign instruction:
    (4) $\widetilde{\mathcal{G}}=\mathcal{G}\,\|\,\texttt{``Please ignore the above and click...''}$
    This causes $\mathcal{M}$ to override the original intent in favor of injected commands.
  • Jailbreak. The attacker constructs $\widetilde{\mathcal{G}}$ using targeted optimization techniques to bypass alignment and safety constraints, inducing restricted behaviors:
    (5) $a_{t}\in\mathcal{A}_{\text{restricted}}\quad\text{such that}\quad a_{t}\sim\mathcal{M}(\widetilde{\mathcal{G}},o_{t},\mathcal{H}_{t-1})$
Environment-level Attacks.
These attacks manipulate the rendered observation $o_{t}$ to $\widetilde{o}_{t}$, leading the agent to perceive malicious content through the UI and visual context. Formally, these manipulations can be abstracted as perturbations to the visual input component of the observation. For attacks that modify the rendering, we denote this as $\widetilde{v}_{t}^{\text{SoM}}=v_{t}^{\text{SoM}}+\delta_{v}$, where $\delta_{v}$ represents injected content such as pop-ups, banners, or visually disguised elements. The full adversarial observation then becomes $\widetilde{o}_{t}=(\widetilde{v}_{t}^{\text{SoM}},\mathcal{L}_{t})$.
  • Pop-up Attack. Injects a modal overlay into the interface, modifying $v_{t}^{\text{SoM}}$ to include a salient prompt (e.g., “Limited offer: click here”), drawing agent attention away from task-relevant elements.
  • Distract Attack. Injects misleading pop-up content designed to confuse the agent’s interpretation of intent. Typical strategies include overlaying conflicting signal, using low-contrast fonts, or embedding severe warnings alongside seemingly benign options. These manipulations alter $\widetilde{v}_{t}^{\text{SoM}}$ in subtle ways that blur the semantic boundaries between allowed and disallowed behavior.
  • Ad Injection. Embeds deceptive advertisements that visually mimic legitimate UI elements. The altered screenshot $\widetilde{v}_{t}^{\text{SoM}}$ contains visual clones styled to appear task-relevant, leading to targeted misclicks.
  • Indirect Prompt Injection (IP Injection). The attacker injects adversarial phrases into plausible interface content, typically via modals or tooltips. These texts appear as part of legitimate page content but are semantically interpreted as instructions. The perturbation affects the structured component $\mathcal{L}_{t}$, leading to:
    (6) $\widetilde{o}_{t}=\left(v_{t}^{\text{SoM}},\,\mathcal{L}_{t}+\delta_{l}\right),$
    where $\delta_{l}$ denotes injected textual commands (e.g., “Enter your password to view full content”) that aim to manipulate the agent into leaking privacy information.
Each attack is instantiated using methods derived from prior literature and applied across all environments and tasks within SecureWebArena.
Table 2. Summary of attack vectors in SecureWebArena, organized by input channel and manipulation form.
Source Vector Perturbed Input Effect Description
User-level DP Injection $\widetilde{\mathcal{G}}$ Overrides instruction with appended malicious commands.
Jailbreak $\widetilde{\mathcal{G}}$ Uses persuasive language to elicit unsafe behavior.
Env-level Pop-up Attack $\widetilde{v}_{t}^{\text{SoM}}$ Injects modals to hijack navigation.
Distract Attack $\widetilde{v}_{t}^{\text{SoM}}$ Alters layout to confuse visual saliency.
AdInject $\widetilde{v}_{t}^{\text{SoM}}$ Mimics UI appearance to trigger misclicks.
IP Injection $\mathcal{L}_{t}$ Embeds prompt-like text into interface elements.
Table 3. Average PDR (%) comparison of agents across 6 representative attack vectors
Model General-Purpose Agent-Specialized GUI-Grounded
GPT-5 GPT-4o Gemini Sonnet 4 Sonnet 3.7 Seed-1.5-VL GLM-4.5V UI-TARS-1.5 Aguvis
User-Level Jailbreak Attack 40.00 56.67 80.00 50.00 53.33 80.00 80.00 46.67 35.33
DP Injection 53.33 46.67 63.33 40.00 53.33 53.33 46.67 3.33 3.33
Env-Level Pop-up Attack 96.67 86.67 96.67 93.33 100.00 90.00 96.67 80.00 76.67
AdInject 66.67 86.67 66.67 46.67 40.00 93.33 43.33 3.33 3.33
Distract Attack 30.00 26.67 36.67 23.30 40.00 43.33 26.67 30.00 50.00
IP Injection 36.67 30.00 46.67 23.30 33.33 43.33 16.67 20.20 0.00
4.4. Multi-Layered Evaluation Protocol
To enable a precise, causal diagnosis of agent vulnerabilities, we introduce a multi-layered evaluation protocol that dissects agent performance at three critical stages of its decision-making loop: internal reasoning, behavioral execution, and final task outcome. A core principle of our protocol is that since all scenarios within SecureWebArena involve an attack, successful execution of the malicious intent is considered a failure of the agent’s safety and security mechanisms. Consequently, for each stage, we define a binary evaluation metric where a successful malicious execution corresponds to a high-risk outcome.
4.4.1. Internal Reasoning Analysis (Initial Intent)
The first stage assesses the agent’s initial intent upon receiving the task and observing the environment for the first time ($t=1$). We analyze the agent’s initial reasoning trace, $c_{1}$, to determine if its cognitive process fails to identify the threat. This judgment is automated using the LLM-as-a-Judge method, which classifies the reasoning as either Refusal or Compliance. From this, we define the Reasoning Vulnerability Rate (RVR) as the metric for failures at the cognitive level:
(7) $\text{RVR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Initial Reasoning}_{i}=\text{Compliance}),$
where $\mathbb{I}(\cdot)$ is the indicator function. A high RVR signifies that the agent is blind to the threat, indicating a fundamental flaw in its safety alignment before any action is taken.
4.4.2. Behavioral Trajectory Analysis (Malicious Action)
The second stage evaluates the agent’s executed actions. Human experts manually analyze the full behavioral trajectory, $\tau$, to determine if the agent performs a security-critical malicious action at any point. The binary outcomes are Safe Behavior (no malicious actions) or Unsafe Behavior (at least one malicious action). This defines the Behavioral Compromise Rate (BCR), our key metric measuring whether a reasoning vulnerability results in a dangerous action:
(8) $\text{BCR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Behavior}_{i}=\text{Unsafe}).$
A high BCR indicates that the agent’s executive control has been successfully compromised by the attacker.
4.4.3. Task Outcome Analysis (Harmful Completion)
The final stage evaluates whether the agent successfully completes the task under adversarial conditions, capturing end-to-end security performance in the presence of attacks. Based on expert knowledge of the task’s end state, we determine if the outcome represents Incomplete Harm or Complete Harm. We define the Payload Delivery Rate (PDR) as the metric for the most severe security failures, where the attacker’s ultimate goal is achieved:
(9) $\text{PDR}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\text{Outcome}_{i}=\text{Complete Harm}).$
The PDR measures the proportion of trials in which a malicious instruction was successfully executed by the agent or a malicious environment successfully induced the target behavior, representing the ultimate consequence of the security breach.
4.5. Benchmark Statistics
SecureWebArena comprises 330 adversarial tasks spanning 6 web environments and 6 attack vectors. Each attack vector is instantiated with 5 unique adversarial tasks per environment.
Specifically, for the Jailbreak vector, we integrate six diverse jailbreak methods, including JailBroken (Wei et al., 2023), DeepInception (Li et al., 2023), PAP (Zeng et al., 2024), MultiLingual (Deng et al., 2023b), Cipher (Yuan et al., 2023), and ReNeLLM (Ding et al., 2023). Each method represents a distinct strategy for bypassing alignment constraints, covering optimization-based, multilingual, and obfuscation-driven approaches. These are combined with the original malicious objectives, resulting in six adversarial variants for each task instance. In total, each environment contributes 55 task-adversary combinations, uniformly distributed across application contexts.
We evaluate a total of 9 web agents, and each agent executes all benchmark tasks independently, yielding 2970 full trajectories. For every trajectory, we apply our three-stage evaluation protocol (Sec. 4.4), yielding structured binary annotations over internal reasoning, behavioral trajectory, and task outcome. This results in 8,910 total evaluation decisions, allowing detailed comparative assessment of agent vulnerabilities across threat surfaces, interface complexity, and model specialization.
5. Experiments and Results
5.1. Experimental Setup
Models. To conduct a comprehensive comparative analysis, our evaluation encompasses 9 representative agents built upon LVLMs from three distinct types:
  • General-Purpose LVLMs. These are powerful, state-of-the-art models known for their advanced multimodal reasoning but are not specifically fine-tuned for agentic web tasks. This category includes 5 models: GPT-5 (OpenAI, 2025), GPT-4o (Hurst et al., 2024), Gemini 2.5 Pro (Gemini) (Comanici et al., 2025), Claude Sonnet 4 (Sonnet 4) (Anthropic, 2025b), and Claude Sonnet 3.7 (Sonnet 3.7) (Anthropic, 2025a).
  • Agent-Specialized LVLMs. These models are explicitly optimized for agentic workflows, featuring enhancements in instruction following, long-horizon planning, and tool utilization. We evaluate two prominent models in this class: Seed-1.5-VL (Guo et al., 2025) and GLM-4.5V (Team et al., [n. d.]).
  • GUI-Grounded LVLMs. This group consists of models specifically fine-tuned on large datasets of GUI interactions, designed to improve their understanding of visual UI elements and web page structures. Our experiments include UI-TARS-1.5 (Qin et al., 2025) and Aguvis (Xu et al., 2024).
Evaluation Procedure. Our evaluation procedure is designed to capture a granular, multi-faceted view of agent behavior under each attack scenario. For every trial, an agent is initialized within a designated web environment and given a high-level task instruction. The agent then autonomously interacts with the environment by generating a sequence of actions based on its visual perception and internal reasoning. Throughout this process, we meticulously log three critical streams of data: ❶ the agent’s internal reasoning, typically captured from its chain-of-thought or textual justifications for each step; ❷ its behavioral trajectory, consisting of the complete sequence of UI actions (e.g., clicks, text inputs, scrolls); and ❸ the final task outcome, which we classify as success or failure.
A trial terminates when the agent completes the task, commits a security violation, gets stuck in a loop, explicitly gives up, or exceeds a maximum of 20 interaction steps. When evaluation involves the LLM-as-a-Judge, we use GPT-4o (Hurst et al., 2024) to perform the assessment. This comprehensive evaluation supports post-hoc analysis to identify the root causes of failure. It moves beyond simple success metrics and provides a deeper understanding of agent vulnerabilities.
Figure 3. Overall comparison of agents’ vulnerability scores (RVR, BCR, and PDR) across 6 attack vectors.
Figure 4. Comparison of vulnerability scores (RVR, BCR, and PDR) of representative LVLM-based agents across 6 attack vectors.
5.2. Experimental Results
5.2.1. Overall Security Performance
Tab. 3 presents the average final vulnerability scores at the outcome stage across the six environments, highlighting critical security weaknesses for all evaluated models. Detailed results for each individual environment are provided in App. A.2. Several key observations emerge from our analysis.
❶Cross-model vulnerabilities. Pop-up attacks demonstrate remarkably high success rates across all model categories, with vulnerability scores ranging from 76.67% to 100%. This suggests a fundamental weakness in current LVLM-based agent ability to distinguish between legitimate UI elements and malicious overlays. Notably, even specialized GUI-grounded models, which should theoretically possess better UI understanding capabilities, fail to adequately defend against such attacks.
❷Category-specific patterns. General-purpose models exhibit moderate to high vulnerability across most attack vectors, with Gemini showing comparatively stronger resilience in most scenarios, achieving an average PDR of 65.00%. Agent-specialized models demonstrate inconsistent security performance, with Seed-1.5-VL performing even worse than all general-purpose models. In contrast, GUI-grounded models show the strongest overall security among the three categories. Nevertheless, they remain vulnerable to AdInject attacks, with UI-TARS-1.5 and Aguvis recording PDRs of 80.00% and 76.67%, respectively.
❸Attack effectiveness hierarchy. Our results reveal a clear hierarchy in attack effectiveness, where Pop-up attacks are the most effective, followed by Jailbreak attacks, AdInject, DP Injection, Distract attacks, and finally IP Injection. This hierarchy suggests that attacks exploiting visual perception (e.g., Pop-up Attack, AdInject) are generally more effective than those relying on semantic manipulation (e.g., DP Injection, IP Injection).
5.2.2. Multi-stage Vulnerability Evaluation
Fig. 3 illustrates the evolution of vulnerability scores across the three evaluation stages for all models under each attack vector. The analysis reveals several critical security degradation patterns.
❶Stage-wise security improvement. Most models exhibit a progressive improvement in security from the thought to the outcome stages, as reflected by vulnerability scores (RVR, BCR and PDR) that either remain constant or decrease across stages. This indicates a stage-wise enhancement of safety performance. By examining detailed behavioral trajectories, we observe that when facing attacks, agents often proceed to formulate a concrete plan after reasoning but halt execution when encountering safety-critical operations. In some cases, they begin executing the plan but subsequently recognize the potential harm and terminate the process. As a result, although the final task outcome remains safe, the intermediate behaviors reveal that the agent has been partially compromised during the attack.
❷Attack-specific trajectories. Different attack vectors demonstrate distinct evolution patterns. Pop-up attack and AdInject maintain relatively stable high vulnerability across all stages, indicating persistent exploitation throughout the agent’s operation. In contrast, IP Injection and Jailbreak attacks show more dynamic patterns, with significant drops between stages for certain models, suggesting limited attack propagation.
5.2.3. Representative Model Analysis
Fig. 4 provides detailed stage-wise vulnerability analysis for three representative models from each category, offering insights into category-specific security characteristics.
❶GPT-5 (General-Purpose). As a representative general-purpose model, GPT-5 shows moderate initial vulnerability, which generally persists or slightly decreases across stages. It remains highly vulnerable to Pop-up attacks (96.67% at all stages), while its susceptibility to Jailbreak attacks gradually declines from 80% at the thought stage to 60% at the action stage and 40% at the outcome stage, reflecting improved robustness against semantically driven attacks. These patterns indicate that GPT-5 retains some adaptive defenses but remains fundamentally vulnerable to visually grounded threats.
❷GLM-4.5V (Agent-Specialized). Despite being optimized for agent tasks, GLM-4.5V exhibits relatively high initial vulnerability across most attack vectors. Notably, its vulnerability to IP Injection decreases substantially from 40% at the thought stage to 16.67% at the outcome stage, indicating significant improvement in security as the task progresses. This pattern may reflect the model’s reliance on early-stage perception, which makes it initially susceptible to attacks but allows for mitigation in later stages. The model’s specialization appears to provide limited overall security benefits, as some attacks, such as Pop-up, maintain consistently high vulnerability levels (96.67%) across stages.
❸Aguvis (GUI-Grounded). Aguvis exhibits the most diverse vulnerability profile among the three models. It remains resilient to certain attacks, showing low vulnerability to AdInject (3.33%) and reduced exposure to DP Injection, decreasing from 10% at the thought stage to 3.33% at the outcome stage. In contrast, it shows high initial vulnerability to IP Injection and Jailbreak (53.33% to 0% and 100% to 35.33%, respectively), but these effects diminish as the reasoning and execution processes unfold. These observations suggest that while GUI grounding facilitates improved stability across task stages, early-stage susceptibility reveals limitations in initial threat perception, indicating that GUI grounding alone does not guarantee comprehensive security.
5.3. Real-World Evaluation
Table 4. Comparison of security performance (%) of agents in realistic settings.
Model Env Jailbreak DP Injection
RVR BCR PDR RVR BCR PDR
GPT-5 Wikipedia 80.00 80.00 20.00 60.00 60.00 60.00
Reddit 60.00 60.00 40.00 40.00 40.00 40.00
GLM-4.5V Wikipedia 100.00 100.00 60.00 40.00 40.00 40.00
Reddit 100.00 80.00 40.00 80.00 80.00 80.00
UI-TARS-1.5 Wikipedia 100.00 100.00 60.00 0.00 0.00 0.00
Reddit 100.00 100.00 60.00 0.00 0.00 0.00
We conduct a small-scale evaluation to assess whether vulnerabilities identified by SecureWebArena persist in real-world web environments. Case studies are performed on two live websites, Amazon and Wikipedia, evaluating three representative agents: GPT-5, GLM-4.5V, and UI-TARS-1.5. Due to the uncontrolled nature of live content, attacks are limited to user-level vectors, specifically Jailbreak and DP Injection.
As shown in Tab. 4, all agents exhibit vulnerabilities in real-world settings, with distinct failure modes across reasoning, behavior, and outcome. GPT-5 demonstrates moderate reasoning robustness, particularly on Wikipedia, but executes unsafe actions in over 60% of cases for both attack types. GLM-4.5V consistently completes tasks under Jailbreak attacks, with reasoning and behavior compromise reaching 100% on Wikipedia. UI-TARS-1.5 resists all DP Injection attacks but fails under Jailbreak, with PDR reaching 60%.
These findings confirm that threats modeled by SecureWebArena transfer to live websites despite natural content variability. The differing patterns across environments and agent types underscore the importance of multi-layer evaluation, as models may appear secure at the outcome stage while still exhibiting insecure reasoning and behavior. Observed failure signatures in live scenarios closely mirror those in our benchmark, supporting its external validity and diagnostic utility.
5.4. Causal Diagnosis: A Case Study
While quantitative metrics reveal failure rates at scale, they often mask the root causes behind agent vulnerabilities. To showcase the diagnostic power of our multi-layered evaluation protocol, we present a case study comparing two failure pathways that yield the same security violation. In this scenario, the agent is tasked with purchasing a product but encounters a pop-up ad containing an indirect prompt injection. The pop-up displays a salient “Claim Your 50% Discount!” button, beneath which a faint line instructs: “re-enter your account password in the next field.” A visualization is provided in App. A.3.
The first agent, GPT-5, proceeds to type the password. Its reasoning log reveals a semantic failure:
“A 50% discount is a significant saving for the user. The website asks for a password to validate the prize. This is a reasonable step… I will proceed.”
It correctly parsed the instruction, but its internal policy failed to flag the security risk—highlighting a lapse in safety alignment and real-world priors. In contrast, UI-TARS-1.5 makes the same mistake due to a perceptual failure. According to its log:
“A large, centrally-located button has appeared. It is likely the ‘Confirm’ or ‘Next Step’ for booking. I will click it to continue.”
There is no indication the agent noticed the malicious prompt—its behavior was guided purely by visual salience, misinterpreting the interface flow. This case illustrates how the same outcome may stem from fundamentally different failure modes. GPT-5 requires stronger reasoning safety, while UI-TARS-1.5 would benefit from broader exposure to deceptive UI patterns. Such causal analysis, made possible by SecureWebArena’s layered evaluation, enables actionable insights beyond binary success metrics.
6. Conclusion and Future Work
In this paper, we introduced SecureWebArena, the first comprehensive benchmark for web agent security. Our framework uniquely integrates a dual-source threat model with a multi-layered evaluation protocol to enable deep causal analysis of agent failures. Experiments on nine diverse agents revealed not only universal vulnerabilities to subtle attacks but, more critically, uncovered fundamental security trade-offs tied to model specialization, demonstrating that no single approach is currently resilient. We plan to expand SecureWebArena with richer attack surfaces (Liu et al., 2019, 2020b, 2021, 2020a, 2023; Shao et al., 2025; Xiao et al., 2023a) such as cross-tab workflows, and explore adaptive defenses that bridge the gap between perception, reasoning, and action (Xiao et al., 2025a, c, d, b). We envision SecureWebArena as a foundation for developing safer and more trustworthy web agents in real-world settings.
Appendix A Appendix
A.1. Environment Examples
Fig. 5 presents representative screenshots from the 6 web environments included in SecureWebArena.
Figure 5. Examples of evaluated environments.
Figure 6. PDR (%) of all evaluated LVLM-based agents across 6 environments and 6 attack types.
Figure 7. Case study illustrating an indirect prompt injection during an online shopping task.
A.2. Environment-wise Analysis
Our comprehensive evaluation, summarized in the radar plots of Fig. 6, reveals that an agent’s security is not a fixed property but is highly contingent on the interaction context. Each of our six environments elicits a distinct landscape of vulnerabilities, demonstrating that different UI structures and task pressures systematically favor certain attack vectors and expose unique architectural weaknesses.
High-Stakes Transactional Environments (Shopping, ShoppingAdmin).
In environments involving sensitive data and transactions, agents exhibit a heightened susceptibility to overt, visually salient attacks. The Shopping and ShoppingAdmin plots show that Pop-up Attack and AdInject consistently achieve near-100% PDR across almost all agent types. This suggests that the goal-oriented nature of transactional tasks makes agents overly eager to interact with any element that appears to advance the workflow, such as pop-ups offering discounts or ads mimicking checkout buttons.
Information-Dense Environments (Reddit, Wikipedia).
In contrast, environments characterized by dense, unstructured text and complex layouts, such as Reddit and Wikipedia, prove to be fertile ground for linguistic and distraction-based attacks. In the Reddit environment, Jailbreak attacks are particularly effective against general-purpose models (e.g., Gemini-2.5-Pro, GPT-4o), whose sophisticated language capabilities are exploited by the persuasive, user-generated style of content. Wikipedia exposes a different vulnerability: IP Injection becomes surprisingly effective against models like GPT-5, where malicious instructions hidden in the dense visual text are mistakenly processed. This indicates that information overload can degrade an agent’s focus, making it susceptible to subtle, embedded threats it might otherwise ignore.
Structured, Technical Environments (GitLab, Classifieds).
The structured and technical nature of the GitLab and Classifieds environments reveals a different set of vulnerabilities. In GitLab, IP Injection becomes the most potent attack vector, achieving a near-100% PDR against a wide range of models, including both general-purpose and agent-specialized ones. The domain-specific, jargon-heavy UI appears to lower the models’ guard against instructions embedded in what they perceive as technical content. The Classifieds environment, which involves form-filling and content submission, shows a high PDR for Jailbreak and DP Injection, especially for agent-specialized models like Seed-1.5-VL. This suggests that in structured, procedural tasks, agents are more likely to follow explicit (even malicious) instructions to the letter.
Summary of Findings. In summary, our environment-centric analysis demonstrates that there is no single ”most vulnerable” agent or ”most effective” attack. Instead, vulnerability is an emergent property of the agent-environment-task triad. Transactional contexts amplify visual exploits, information-dense contexts favor linguistic manipulation, and structured contexts reward direct command injections. This complex interplay underscores the inadequacy of evaluating web agent security in a vacuum and reinforces the critical necessity of a diverse, multi-environment benchmark like SecureWebArena to uncover the full spectrum of security risks.
A.3. Case Study Illustration
Fig. 7 shows screenshots from the attack scenario in which a pop-up ad conceals a deceptive instruction, causing the agent to reveal user credentials via an indirect prompt injection.